# Solvent: liquidity verification of smart contracts

### v2

@@ 0.A.3 Related work
Besides Certora, only SmartPulse [31] and VeriSolid [25] seem capable to deal with liveness properties. SmartPulse targets directly Solidity code, and it has a property specification language based on Linear Temporal Logic (LTL). This makes it possible to express liveness properties, but not liquidity properties, that are out of the reach of LTL (see Section 0.A.2). Therefore, the properties supported by SmartPulse and Solvent have uncomparable expressiveness. To be usable in practice, liveness properties must be accompanied by a fairness assumption, i.e. another LTL formula that specifies the traces where a user performs some required action in order to reach the desired state. For example, in our crowdfunding contract (Section 2) a target property could be “if the target is not reached, then a donor gets its money back”, and the associated fairness assumption could be “the donor performs the action wdDonor”. A main difference between Solvent and SmartPulse is that, while in SmartPulse the designer of the property must anticipate, in the fairness assumption, the sequence of transactions that a user must perform in order to reach a certain state change, in Solvent this sequence is inferred by the the SMT solver. In particular, the solver infers the (minimal) length of the sequence, the called methods, and their actual parameters in order to produce the desired state change.
@@
0.A.4 Use case: a 2-players lottery

### v3

@@ 0.A.2 Verification of liquidity in other tools
Besides Certora, only SmartPulse [34] and VeriSolid [28] seem capable to deal with liveness properties. SmartPulse targets directly Solidity code, and it has a property specification language based on Linear Temporal Logic (LTL). This makes it possible to express liveness properties, but not liquidity properties, that are out of the reach of LTL (see Section 2). Therefore, the properties supported by SmartPulse and Solvent have uncomparable expressiveness. To be usable in practice, liveness properties must be accompanied by a fairness assumption, i.e. another LTL formula that specifies the traces where a user performs some required action in order to reach the desired state. For example, in our crowdfunding contract (Section 3) a target property could be “if the target is not reached, then a donor gets their money back”, and the associated fairness assumption could be “the donor performs the action wdDonor”. A main difference between Solvent and SmartPulse is that, while in SmartPulse the designer of the property must anticipate, in the fairness assumption, the sequence of transactions that a user must perform in order to reach a certain state change, in Solvent this sequence is inferred by the the SMT solver. In particular, the solver infers the (minimal) length of the sequence, the called methods, and their actual parameters in order to produce the desired state change.
@@
0.A.3 Use case: a 2-players lottery